OpenStudy (anonymous):

HOW TO GRAPH: –3x + 2y = 8

sam (.sam.):

–3x + 2y = 8 2y =3x+ 8 y=3x/2+4 y-intersect at 4

sam (.sam.):

You need to make the equation into standard y=mx+c form then you only know the y-intersect

sam (.sam.):

Then you make a table _____________ x | y | ------------ 1 11/2 0 4 Then plot the points